# LibriQuote: A Speech Dataset of Fictional Character Utterances for Expressive Zero-Shot Speech Synthesis

### v1

@@ Introduction
LLMs have shown promising results in literary dialogue and point of view understanding (Piper and Bagga 2024; Michel et al. 2025). Our experiments involved prompting Phi-4888https://huggingface.co/microsoft/phi-4 to extract narrative information for a target quotation in a contextual window of one paragraph before and after the quotation. We use few-shot prompting with 5 examples. Prompt details can be found in the Technical Appendix. To mitigate extraction errors and keep only informative narrative content, we replace all quotations in context with special markers, such that only narration and structure remains. Besides, we experimented with prompting the LLM to self-report a confidence score on a 1-to-10 scale that we used to prune-out unconfident predictions in order to maximize precision (Phi-4).
@@
In an initial phase developed to setup guidelines and discussing complex cases, two annotators processed independently 100 quotations. After discussion and once an agreement was reached, a single annotator continued 300 extra instances, resulting in 400 quotations annotated with verbs, adverbs, nouns and adjectives. We provide in Table 2 the results for verbs and adverbs with Phi-4. Results for other LLMs, nouns and adjectives and runtimes can be found in the Technical Appendix. We see that Phi-4 performs relatively well at extracting the narrative information, but still suffers from hallucinations – in particular for adverbs – as shown by the precision score. When pruning out all predictions with a confidence score lower than 10, Phi-4 excels in terms of precision at the cost of lower recall for both verbs and adverbs. A high precision is desirable: we want to avoid quotations to be wrongly affected with expressive speech verbs and adverbs.
Filtering
@@
We extract narrative information on LibriQuote-train using Phi-4, yielding a (potentially empty) set of predicted verbs and adverbs for each quotation. To produce a filtered subset , we first include all quotations that had a non-empty adverb pseudo-label regardless of the predicted speech verb, as adverbs often give precise information of how a quotation is uttered (e.g. “he said cautiously”). Then, based on a manually defined set of expressive speech verbs999We kept 89 out of 201 speech verbs from the following list: https://archiewahwah.wordpress.com/speech-verbs-list/, we include all quotations (which were not kept in the first step) that have a verb falling into that list. As a result, , contains 377,776 quotations (11% of the full quotation train set) for a total of 379 hours. We believe can be used for data-efficiency, or can foster analysis on the alignment between written narrative information and uttered speech.
Experiments
In this section, we describe our main LibriQuote experiments and further evaluations on benchmark datasets.
LibriQuote Test Set
Baselines
To evaluate to which extent state-of-the-art TTS systems are able to synthetize expressive speech, we compare three TTS systems: SparkTTS (Wang et al. 2025), F5-TTS (Chen et al. 2024) and MaskGCT (Wang et al. 2024). These models were chosen based on their availability and performance on standard benchmarks. SparkTTS is an autoregressive TTS model based on Qwen2-0.5B. It generates global and semantic discrete tokens derived from BiCodec (Wang et al. 2025), that are then converted back to raw audio. Global tokens capture speaker information such as timbre, while semantic tokens capture fine-grained semantic information. F5-TTS is a non-autoregressive flow matching TTS system based on Diffusion Transformer (Peebles and Xie 2023a). MaskGCT is a non-autoregressive TTS system, that follows a mask-and-predict paradigm. It follows a two-stage process: semantic tokens derived from a VQ-VAE model with Wav2Vec-Bert 2.0 features are predicted in a first step, and then used to condition the prediction of acoustic tokens. They were all trained with approximately 100K hours of speech data. We use publicly available checkpoints for each model and do not modify generation parameters. Baselines details are available in the Technical Appendix.
| LibriSpeech-PC | SeedTTS test-en | |||
@@
| F5-TTS | 2.42 | 0.66 | 1.83 | 0.67 |
| MaskGCT | - | - | 2.62 | 0.72 |
| SparkTTS | 2.10⋆ | 0.51 | 2.07 | 0.42⋆ |
| SparkTTS | 2.00⋆ | 0.51 | 1.90⋆ | 0.42⋆ |
Metrics
Our evaluation follows the cross-sentence design (Le et al. 2023): a narration utterance of 2 to 15 seconds is used as audio context to guide the generation of a quotation utterance. For objective metrics, we report Word Error Rate (WER) computed with Whisper-large-v3 (Radford et al. 2023) to measure speech intelligibility and speaker similarity between the synthesized speech and the original ground-truth speech (SIM-O). We employ a WavLM-large based speaker verification model (Chen et al. 2022) to extract speaker embeddings and calculate cosine similarities. We evaluate prosody similarity with the ground truth quotation using Mel Cepstral Distortion (MCD) and F0 Pearson correlation (FPC) following (Huang et al. 2023). We also evaluate emotion and accent similarity (E-Sim, A-Sim) by leveraging representations from dedicated pretrained models: Emotion2Vec (Ma et al. 2024) and an ECAPA-TDNN model trained on Common Voice (Desplanques, Thienpondt, and Demuynck 2020; Ardila et al. 2020; Zuluaga-Gomez et al. 2023). Similarity scores are calculated by computing the cosine similarity between the ground-truth and synthesized speech representations. For subjective metrics, we employ Mean Opinion Score (MOS, rated from 1 to 5 with 0.5 intervals) to assess naturalness, and Comparative MOS (CMOS, rated from -3 to 3) to measure the degree of expressivity with respect to the ground-truth. We randomly selected 2 samples per speaker in LibriQuote-test, for a total of 30 samples. Each sample was judged by 5 raters. More details can be found in the Technical Appendix.
Analysis
As shown in Table 3, the ground-truth WER is relatively high compared to standard audiobooks datasets such as LibriSpeech. We believe this high WER is likely due to an higher expressiveness in our test samples. SparkTTS achieves the lowest WER (5.5) across models, but the lowest SIM-O. In contrast, MaskGCT achieves the highest WER and SIM-O. F5-TTS and MackGCT synthesize speech with higher accent and emotion similarity than SparkTTS. However, in terms of prosody similarity, SparkTTS performs better, with an higher FPC and lower MCD. The subjective evaluation reveals that all models fail to synthesize speech as expressive as the ground-truth utterances: the latter were consistently preferred over synthesized speech in the CMOS experiment. In addition, ground-truth quotations appear more natural than generated speech, with a MOS of 3.58, while F5-TTS achieves the best MOS of 3.47. These results suggest that the cross-sentence task of LibriQuote-test is challenging, even for modern TTS systems.
Fine-Tuning Experiments
Experiments
We evaluate how and can be used to enhance TTS systems by fine-tuning SparkTTS on these corpus. We chose SparkTTS as a baseline for its simplicity of implementation, which does not requires flow-matching as F5-TTS and MaskGCT.
We compute global and semantic tokens for each training utterances with BiCodec. Then, we fine-tune SparkTTS using the standard language modeling task on semantic tokens with text and global tokens prepended on the two corpora (SparkTTS and SparkTTS). We employ Adam optimizer (Kingma and Ba 2017) with peak learning rate of 1e-5. We train each model for 3 epochs on their respective data. A cosine schedule is used with 10000 warmup steps for SparkTTS and 5000 steps for SparkTTS (note that SparkTTS is only trained for 10% of the total steps of SparkTTS), and we form batches of 32 utterances. During inference, we compute global tokens using reference narration utterances, and prepend the text to synthesize and global tokens before starting the generation.
Datasets & Metrics
We use LibriSpeech-PC (Meister et al. 2023), a punctuation restored version of LibriSpeech – our training data also contains punctuations – with the standard split from Chen et al. (2024) that contains 1127 samples with 4-to-10 seconds audio prompts. We also evaluate on SeedTTS test-en (Anastassiou et al. 2024), which is composed of 1088 samples from Common Voice (Ardila et al. 2020). We do not report scores for the Mandarin variant of SeedTTS as our fine-tuning data only support English speech, although it would be interesting to see how fine-tuning impacts performance on other languages. We report WER computed with Whisper-large-v3 (Radford et al. 2023) and speaker similarity to the reference speech (SIM), computed with WavLM-large (Chen et al. 2022). We compare to SparkTTS, F5-TTS and MaskGCT (only SeedTTS test-en scores are reported for the latter). We report scores computed with samples we generated for SparkTTS on SeedTTS test-en as we were not able to reproduce the original scores.
Analysis
Results are displayed in Table 4. We see that SparkTTS achieves the lowest WER on LibriSpeech-PC, with a large improvement over standard SparkTTS. This gain is also seen on SeedTTS test-en, at the cost of a decrease in SIM. Interestingly, when fine-tuning on the filtered subset , we see a large improvement over SparkTTS, where SparkTTS almost reaches SparkTTS WER and similar SIM, with only 10% of the fine-tuning data and training steps. Thus, reveals as a data-efficient alternative to the full training set for speech intelligibility. These results suggest that training with LibriQuote-train provides substantial improvement to speech intelligibility, while maintaining speaker similarity on audiobook data.
Training from Scratch
| WER | SIM-O | FPC | MCD | |
| SparkTTS | 5.5 | 0.453 | 0.65 | 9.1 |
| SparkTTS | 13.3⋆ | 0.394⋆ | 0.64 | 10.1⋆ |
| context | 11.3⋆ | 0.396⋆ | 0.65 | 9.7⋆ |
| SparkTTS | 6.5 | 0.465⋆ | 0.66⋆ | 8.9⋆ |
| SparkTTS | 10.3⋆ | 0.458⋆ | 0.66 | 9.1 |
Additionaly, we train two variants of SparkTTS from scratch using the quotation subset . We initialize SparkTTS with Qwen2-0.5B weights, and continue training by learning to autoregressively predict the next semantic tokens, conditioned on the text and global tokens. Additionally, we leverage the contextual information around quotation by replacing the text condition with the sequence where and are the text tokens from the left and right context respectively, are the tokens from the text to synthesize, and and are special delimiter tokens. Note that we replace each quotation appearing in and with a special token. Therefore, most of the contextual information can be seen as narrative details. We use a context of one paragraph for both and . We train each variant for 10 epochs with an Adam optimizer, a peak learning rate of 3e-5 and a batch size of 32. A cosine schedule is used with 20000 warmup steps.
Analysis
Results on LibriQuote-test are displayed in Table 5. The model trained from scratch typically suffers from larger WER and lower SIM-O than the original SparkTTS. This is expected, as only 5K hours of training data were used, which is small compared to the 100K hours used by SparkTTS. The same trend follows for FPC and MCD, indicating lower prosody similarity to the ground-truth. A typical problem encountered during generation is word hallucination, which we believe is likely due to the insufficient amount of data to train such autoregressive models. Interestingly, when using contextual information, we see an improvement in WER, FPC and MCD, indicating better speech intelligibility and prosody, which is likely due to the extra narrative details provided. Designing systems capable of leveraging such contextual information might be key to improve the expressiveness of current TTS model.
We also report scores of fine-tuned models in the table: SparkTTS yield a larger WER, but improve slightly the speaker similarity over SparkTTS, while SparkTTS improves over all metric.
Conclusion
This work introduces LibriQuote, a publicly available speech dataset derived from audiobooks, designed for training and benchmarking expressive zero-shot TTS systems. The LibriQuote training set contains 12,720 hours of neutral, non-expressive narration utterances and 5359 hours of more expressive speech in the form of quotations from fictional characters in books, while the test set comprises 7.5 hours of quotations from 15 speakers. Our qualitative analysis reveal that LibriQuote-test covers a large set of emotions and accents. Additionally, we supplement each utterance with the book context in which it was written and provide pseudo-labels for the speech verbs and adverbs used to introduce each quotation. Objective and subjective evaluations show that the TTS systems we evaluated fail to synthesize speech as expressive as the ground-truth utterances and that LibriQuote-train improves speech intelligibility when used as a fine-tuning corpus. We believe that the release of this dataset can foster research on expressive TTS and on contextual conditioning for audiobook synthesis.
Limitations
This work bears several limitations. First, we did not filter LibriQuote utterances that exhibit very high WER between the transcript and ground-truth text as done in previous work. Thus, LibriQuote might contain a small number of outliers, which can hinder learning of ASR or TTS systems. We plan to perform WER-based filtering in a future dataset release. Second, our objective analysis of expressiveness remains limited, with metrics such as FPC and MCD that assume a ground-truth sample is deterministic, hence penalizing diversity in generation. We believe subjective listening tests are more suited for evaluation of expressiveness by providing human ratings. Third, our training experiments were solely conducted with SparkTTS, an autoregressive TTS system which bears its own limitations (such as low speaker similarity), and did not involve other system types. However, we believe that these baseline system results are still a valuable source of information. Finally, we experimented with narrative context conditioning in utterance synthesis, but we believe further research should be conducted on how TTS systems may benefit from it.
Ethical Considerations

### v2

@@ 3.3 Contextual Cues for Direct Speech
We use few-shot prompting with 5 examples. To mitigate extraction errors and keep only informative narrative content, we replace all quotations in context with special markers, such that only narration and structure remains. Our experiments involved prompting to self-report a confidence score on a 1-to-10 scale that we used to prune-out unconfident predictions in order to maximize precision. More details are provided in Appendix B.
@@
In an initial phase dedicated to developing guidelines and solving complex cases, two annotators independently tagged 100 quotations. After discussion and once satisfactory agreement was reached (we report a Cohen’s score of 0.87 for verbs and adverbs), a single annotator continued with 300 extra instances, yielding a total of 400 quotations annotated with verbs, adverbs, nouns, and adjectives, drawn from the train split. We found that adjectives and nouns were relatively rare in the annotated data, and that LLMs failed entirely to extract them. We thus report results only for verbs and adverbs in Figure 2.
We see that most models perform relatively well at extracting narrative cues, but still suffers from hallucinations, in particular for adverbs. When pruning out all predictions with a confidence score lower than 10, all models exhibit higher precision at the cost of lower recall for both verbs and adverbs. A high precision is desirable: we want to avoid quotations to be tagged with wrong expressive speech verbs and adverbs. We thus select Phi-4 as the base LLM for further extraction of contextual cues on the full corpora, as it obtained the highest precision score on adverbs and a satisfactory precision on verbs, while also being the smallest and fastest model.
High-Expressivity Split
@@
After extracting contextual cues on the full dataset with the strategy described above, we build a high-expressivity subset. First, we include all quotations that had a non-empty adverb pseudo-label regardless of the predicted speech verb, as adverbs often give precise information on how a quotation is uttered (e.g. “he said cautiously”). Then, based on a manually defined set of expressive speech verbs666The full list of verbs is available in Appendix B.2., we add all quotations that have a verb falling into that list. The resulting split, , contains 377,776 quotations (11% of the full quotation train set) for a total of 379 hours, which may be used for data-efficient expressive TTS, or can foster automatic analysis of the alignment between written narrative cues and uttered speech.
| WER | SIM-O | E-Sim | CtxMOS | WR | |
| GT | 6.5 | - | - | 3.55 | - |
| SparkTTS | 4.8 | 0.46 | 0.69 | 2.94 | 38 % |
| FT () | 4.6 | 0.47⋆ | 0.71 | 2.97 | 41% |
| FT () | 5.9⋆ | 0.46⋆ | 0.71 | 2.89 | 37% |
| Scratch () | 9.5⋆ | 0.40⋆ | 0.71 | 3.09⋆⋆ | 38% |
| Ctxt | 8.2⋆ | 0.40⋆ | 0.71 | 3.15 | 35% |
| Full () | 5.1⋆ | 0.41⋆ | 0.71 | 3.30 | 37% |
| + FT () | 5.1⋆ | 0.41⋆ | 0.71 | 3.30 | 41% |
| F5-TTS | 6.9 | 0.53 | 0.71 | 2.95 | 31 % |
| FT () | 6.6⋆ | 0.54⋆ | 0.71 | 3.33⋆ | 26 % |
4 Training Experiments
In this section, we analyse the impact of using LibriQuote as a resource for fine-tuning or training TTS systems from scratch. Our goal is to understand 1) how LibriQuote could be used as an alternative to standard TTS datasets such as Emilia He et al. (2024), even though it contains 20 less data and 2) how it can be used to enhance expressivity of pre-trained TTS systems when used as a fine-tuning resource. We mainly conduct experiments with SparkTTS and F5-TTS, that we chose as baselines for better reproducibility. SparkTTS is an autoregressive TTS model based on Qwen2-0.5B, that generates global and semantic discrete audio tokens that can be converted back to raw audio. It was trained on approximately 100K hours of speech data, containing in-the-wild and audiobook corpora. F5-TTS is a non-autoregressive flow-matching based TTS system, trained on 100K hours of speech data. We describe SparkTTS and F5-TTS in details in Appendix F. For each models, we use the available checkpoints and do not modify generation parameters. For SparkTTS, we compute global and semantic tokens for each training utterances, and fine-tune the LLM-backbone using the standard language modeling task on semantic tokens with text and global tokens prepended in the sequence. For F5-TTS, we use the provided fine-tuning script available in the official repository777https://github.com/swivid/f5-tts. Training details including optimizer and hyperparameters can be found in Appendix C and Appendix D.
We design training experiments in two ways: fine-tuning experiments with varying quotation dataset (either the full set of quotations or the expressive subset ), and training from scratch experiments with either the full quotation subset or the full training dataset (both narrations and quotations ). Besides, we experiment with replacing SparkTTS text-condition with contextual information in the vicinity of a target quotation (denoted as Ctxt). By having access to more contextual information, we expect the second variant to improve expressivity over standard SparkTTS. More details can be found in Appendix C.
| LibriSpeech PC | SeedTTS test-en | |||
@@ 4 Training Experiments
| FT () | 2.10⋆ | 0.51 | 2.07 | 0.42⋆ |
| FT () | 2.00⋆ | 0.51 | 1.90⋆ | 0.42⋆ |
| Scratch () | 3.27 | 0.47⋆ | 5.14⋆ | 0.41⋆ |
| Full () | 2.22⋆ | 0.49⋆ | 4.05⋆ | 0.41⋆ |
| + FT () | 2.49⋆ | 0.49⋆ | 4.23⋆ | 0.42⋆ |
| F5-TTS | 2.11 | 0.66 | 1.69 | 0.67 |
| FT () | 2.13 | 0.66 | 1.64 | 0.67 |
4.1 Metrics
Our evaluation follows the cross-sentence design Le et al. (2023): a narration utterance of 2 to 15 seconds is used as audio context to guide the generation of a quotation utterance.
Objective Metrics
We report WER computed with Whisper-large-v3 Radford et al. (2023) to measure speech intelligibility and speaker similarity between the synthesized speech and the original ground-truth speech (SIM-O). We employ a WavLM-large based speaker verification model Chen et al. (2022) to extract speaker embeddings and calculate cosine similarities. We evaluate emotion similarity (E-Sim) by leveraging representations from Emotion2Vec Ma et al. (2024) (emotion2vec-plus-base). Similarity scores are calculated by computing the cosine similarity between the ground-truth and synthesized speech representations.
| WER | SIM-O | E-Sim | ContextMOS | Win-Rate | CMOS | MOS | |
| Ground Truth | 6.5 | - | 3.55 | - | 0.0 | 3.58 | |
| SparkTTS | 4.8 | 0.46 | 0.69 | 2.94 | 38% | -0.97 | 3.26 |
| F5-TTS | 6.9 | 0.53 | 0.71 | 2.95 | 31% | -0.98 | 3.47 |
| MaskGCT | 7.6 | 0.56 | 0.72 | 2.94 | 28% | -0.98 | 3.36 |
| IndexTTS2 | 5.2 | 0.49 | 0.63 | 3.33 | 46% | 0.25 | 3.47 |
| IndexTTS2-Context | 5.4 | 0.50 | 0.64 | 3.45 | 54% | 0.03 | 3.63 |
Contextual Metrics
To measure to which extent synthesized speech matches the quotation context, we leverage recent advances in Large-Audio-Language-Models (LALMs) and use an LALM-as a-Judge approach Manku et al. (2025); Ji et al. (2025). We follow Manku et al. (2025) and use Gemini-2.5 Pro as the Judge model, as it was proved to correlate strongly with human judgements. We build an expressive test-subset containing 201 quotations from LibriQuote-test, sampling only quotations that are matched with a predicted adverb, following the methodology described in Section 3.3. We evaluate synthesized and ground-truth utterances in two ways: a single sample Context Mean Opinion Score (ContextMOS) and a two-sample comparative Win-Rate score as defined in Manku et al. (2025). ContextMOS rates on 1-5 scale how an utterance speaking style matches the book context, while Win-Rate calculates the percentage of times a synthesized utterance delivers a more appropriate speaking style than a ground-truth sample, based on the book context. Details can be found in Appendix E.
4.2 Benchmark Datasets
We leverage standard benchmark datasets to evaluate in-domain and out-of-domain TTS performance. We use LibriSpeech-PC Meister et al. (2023), a punctuation restored version of LibriSpeech with the standard split from Chen et al. (2024) that contains 1127 samples with 4-to-10 seconds audio prompts and SeedTTS test-en Anastassiou et al. (2024), which is composed of 1088 samples from Common Voice Ardila et al. (2020). We report WER and cosine similarity between synthesized speech and reference sample (SIM) with the same models as described above.
4.3 Analysis
Table 2 displays results on LibriQuote-test. The ground-truth WER is relatively high compared to standard audiobook datasets such as LibriSpeech. As discussed in Section 3.2, this is partly due to strong emotions or accents in some samples, which increases recognition errors. In contrast, SparkTTS and its variants fine-tuned with and achieve lower WER than the ground truth. These fine-tuned models show slightly higher SIM-O and E-Sim, while ContextMOS and win rates remain similar, suggesting that fine-tuning on LibriQuote modestly improves intelligibility and speaker similarity but not contextual expressiveness. Conversely, models trained from scratch show the opposite trend: using standard text conditions yields a ContextMOS of 3.09 when using the quotation subset as training data, which increases to 3.15 when book context is used as input. This improved expressivity comes with reduced intelligibility and speaker similarity, likely due to a limited amount of training data. This decrease in intelligibility is drastically reduced when using the both narrations and quotations ( when training from scratch. This model achieves greater expressivity compared to all SparkTTS baselines with a ContextMOS of 3.30. Interestingly, fine-tuning this model on a second-stage with the expressive subset does not yield significant gains, echoing the other fine-tuning experiments.
Compared to SparkTTS, we found that fine-tuning conducted with F5-TTS is able to yield large expressivity gain. When fine-tuning with the filtered quotation subset , F5-TTS achieves better speech intelligibility and improved expressivity, with ContextMOS reaching 0.4 points higher. This notable difference is likely related to the different training objectives: while SparkTTS uses an autoregressive loss, it seems that the flow-matching loss of F5-TTS is better able to pick the important expressive information present in LibriQuote’s quotations.
Benchmark Datasets
In Table 3, we present results for benchmark datasets. We omit the from-scratch model with context because these datasets lack additional contextual information. SparkTTS results are reported using samples generated by our synthesis pipeline to ensure fair comparison. Unlike LibriQuote-test, fine-tuned models achieve lower WER than SparkTTS on LibriSpeech-PC and SeedTTS-test-en, indicating better intelligibility. We attribute this to fine-tuning acting as a data curriculum, where high-quality data is introduced late in training, a strategy used in recent TTS training strategies Atamanenko et al. (2025). On LibriSpeech-PC, SIM is comparable between SparkTTS and FT models, but degrades on the out-of-domain SeedTTS-test-en. The Scratch model shows similar WER to SparkTTS on LibriSpeech-PC but performs much worse on SeedTTS, as expected given the mismatch between audiobook and in-the-wild speech. The SparkTTS model trained on the full dataset ( largely improves on speech intelligibility and similarity over the other from-scratch variant, but still struggles on out-of-domain data. Further fine-tuning this model on the expressive subset slightly decreases speech intelligibility on both datasets, further suggesting that fine-tuning is not necessary for a model already trained on the full LibriQuote dataset.
These experiments conclude that supervised fine-tuning on LibriQuote significantly enhances SparkTTS intelligibility on in-domain and out-of-domain data, and that training from scratch enhances expressivity at the cost of less intelligible speech. Besides, fine-tuning a different TTS backbone yields a different picture, showing improves expressivity and intelligibility for F5-TTS. This difference highlights important research directions in understanding the reasons behind less performant SparkTTS fine-tuned models.
Thus, LibriQuote appears as a promising resource to fine-tune flow-matching based TTS systems to be more expressive, or to train from-scratch autoregressive TTS models for an enhanced expressivity.
5 Benchmarking with LibriQuote-test
In this section, we evaluate the ability of state-of-the-art TTS systems to synthesise speech as expressive as amateur audiobook readers from LibriQuote-test. We focus our evaluation on aspects of speech naturalness and emerging capacities to systematically deliver speech utterances that satisfy an implicit or explicit speech intent (e.g. “whispering” or “happiness”). We compare four recent open-source TTS systems: SparkTTS Wang et al. (2025), F5-TTS Chen et al. (2024), MaskGCT Wang et al. (2024) and IndexTTS2 Zhou et al. (2025). These models were chosen based on their availability and performance on standard benchmarks, and are described in Appendix F. Notably, IndexTTS2 first predicts a 7-class emotion distribution based on an emotion prompt (with the text to synthesize as default), that conditions the synthesized speech. This decoupled approach is well-suited for LibriQuote, as the emotion prediction can be done on contextual information around quotations rather than quotation-text alone. We denote this approach as IndexTTS2-Context. All models were trained with the Emilia dataset He et al. (2024) on approximately 50K to 100K hours of speech data, but SparkTTS and IndexTTS2 also use additional Audiobook speech data. We use publicly available checkpoints for each model and do not modify generation parameters.
5.1 Metrics
We use the same set of objective and contextual metrics defined in Section 4.1. In addition, we conduct a subjective rating with human raters. We employ Mean Opinion Score (MOS, rated from 1 to 5 with 0.5 intervals) to assess naturalness, and Comparative MOS (CMOS, rated from -3 to 3) to measure the degree of expressivity with respect to the ground-truth. We randomly selected 2 samples per speaker in LibriQuote-test, for a total of 30 samples. Each sample was judged by 5 raters. More details can be found in Appendix G.
5.2 Analysis
Results are displayed in Table 4. SparkTTS has the lowest WER (4.8) but the lowest SIM-O, while MaskGCT shows the highest WER and SIM-O. F5-TTS and MackGCT produce speech with relatively high emotion similarity. Both IndexTTS2 variants achieve low WER and high ContextMOS and Win-Rates, with IndexTTS2-Context matching ground truth in Win-Rate (54%). Subjective tests further show IndexTTS2 reaches positive CMOS, meaning it is rated more expressive than ground truth on average. Other TTS systems sound natural (high MOS) but fail to capture appropriate prosody, as indicated by negative CMOS and low ContextMOS, highlighting IndexTTS2’s clear advantage in expressive speech synthesis.
While IndexTTS2 already captures emotional expressivity well, replacing text conditioning with contextual conditioning for emotion prediction further improves contextual metrics, achieving human-level Win-Rate (50%) and MOS naturalness. These results show that decoupling emotion prediction from speech synthesis significantly enhances contextual expressivity. Although IndexTTS2 generates more expressive speech than other baselines, it also yields the lowest E-Sim, likely due to emotion prediction errors from text alone. Incorrect emotion predictions can produce speech that contradicts the intended style, which might mislead listeners in applications such as Audiobook reading. As detailed in Appendix H, Quotations read by amateur speakers do not always convey sufficient emotion, but still render the appropriate style more often than all models. In contrast, IndexTTS2 exhibits a higher proportion of both appropriate and mismatched speaking styles than other models, an issue partially mitigated by IndexTTS2-Context.
6 Conclusion
We proposed LibriQuote, a public speech dataset derived from audiobooks, that includes 12,720 hours of neutral narration and 5,359 hours of expressive quotations from fictional characters read by amateur readers. Our qualitative analysis reveals that LibriQuote-test is more emotionally diverse than previous audiobook corpora. Besides, each utterance of LibriQuote is supplemented with its book context and pseudo-labels of quotation intent, showing a wide-range of intended speaking style. Fine-tuning experiments show both training subsets improve SparkTTS intelligibility on other corpora while training from scratch on the full quotation set enhances expressiveness, and that fine-tuning F5-TTS on LibriQuote’s expressive subset yields substantial expressivity gains. Evaluation on LibriQuote-test shows that IndexTTS2 matches ground-truth expressivity, though some synthesized quotations convey unintended emotions. In-depth analysis revealed that amateur readers sometimes convey insufficient emotion in their rendering of quotations, but TTS systems do so at an even higher proportion. Despite recent advancements, TTS systems require substantial improvements to meet standards of professional audiobook narration.
We also want to highlight LibriQuote’s value for practitioners of Digital Humanities. LibriQuote provides 3 million character quotations in text and speech modalities, from 2,991 public domain novels, with pseudo-labels of quotation intent. This large-scale data offers new perspectives to analyse Audiobook prosody, e.g. comparing how male/female speakers interpret male/female characters Pethe et al. (2025) or understanding stylistic differences of words uttered by characters Michel et al. (2024).
7 Limitations
This work bears several limitations. First, we did not filter LibriQuote-train utterances that exhibit very high WER between the transcript and ground-truth text as done in previous work. Thus, LibriQuote might contain a small number of outliers, which can hinder learning of ASR or TTS systems. We plan to perform WER-based filtering in a future dataset release.
Second, our evaluation of expressivity relies on an LALM-as-a-judge approach, which is inherently biased by the LALM understanding abilities and own cultural biases. While prior works have incorporated such approaches in their evaluation pipelines, we believe these automated results should mostly be used to compare models against each other rather than make individual conclusions about models. Our findings, however, reveal closely aligned trends between human and model judgments, indicating that IndexTTS2 is a substantially more capable TTS system for producing expressive speech.
In addition, although our experiments focused directly on end-to-end TTS, we still note that LibriQuote ; and in particular its filtered subset ; might be well-suited for the training of Neural Audio Codecs. Indeed, numerous Audio Codecs targeting English language use LibriSpeech or LibriTTS as training data Wang et al. (2025); Zhang et al. (2024). We thus believe that exploring the use of LibriQuote as training data for Neural Audio Codecs offers an interesting area of research for higher-fidelity reconstruction of expressive speech.
Finally, we experimented with narrative context conditioning in utterance synthesis, but we believe further research should be conducted on how TTS systems may benefit from it.
8 Ethical Considerations